OpenStudy (anonymous):

Find the error in the following solution, and complete the solution correctly. 6/10=x/12 6·x=10·12 6·x=120 x=120/6 x=20 Can u please explain all ur work how u got the solution

OpenStudy (amistre64):

first step is bad

OpenStudy (amistre64):

it looks like they tried to cross multiply .... another misnomer up there with foil

OpenStudy (amistre64):

mutiply both sides by 10(12) clear the fractions

OpenStudy (amistre64):

\[\frac{6(10)(12)}{10}=\frac{x(10)(12)}{12}\iff6(12)=10x\]

OpenStudy (amistre64):

\[\frac{72}{10} = \frac{10x}{10}\iff\frac{72}{10}=x\]

OpenStudy (anonymous):

Thx! for the help